What is the formula weights of the following? Nal
Answer #1

NaI = (22.98 + 126.90) = 149.88 amu

CCl4 = [12.01 + 4 (35.4530] = 153.82 amu

Na3PO4 = [3(22.98) + 30.97 + 4(15.9994)] = 163.90

UF6 = [238.028 + 6 (18.998)] = 352.016